OpenStudy (anonymous):

The average score on a standardized test is 750 points with a standard deviation of 50 points. If 2,000 students take the test at a local school, how many students do you expect to score between 700 and 800 points? 1130 1200 1360 1500

OpenStudy (anonymous):

P (700 < x < 800) = P (-1 < z < 1) = 0.6827 0.6827*2000 = 1365.4 so it would be 1360

OpenStudy (anonymous):

I have one last one: Suppose you have a mean standardized score of 1500 points with a standard deviation of 150 points. This data is normally distributed. What is the z-score of 1300 points? –0.65 –1.33 –1.50 –1.56

OpenStudy (anonymous):

1300 - 1500 ------------- = -200/150 = -1.33 . 150 A. –0.65 B. –1.33 <==ANSWER C. –1.50 D. –1.56